A student's grade in a course is the average of 4 test grades and a final exam that is worth three times as much as each test. Suppose a student has test grades of 91, 90, 83, and 93. Write an equation to model this situation where x is the student's grade on the final exam and y is the student's average for the course.
Then find the score they will need to receive on their final exam if they want to have a grade of 90 for course.
Final exam score needed =

Answer by JoelSchwartz(130)

x=score on the final exam
y=average score for the course
The number called the divisor is seven. The reason the divisor is seven is there are four tests plus the final exam, which is worth three tests.
(91+90+83+93+x)/7=y
(357+x)/7=y
357/7+x/7=90
51+x/7=90
x/7=90-51
x/7=39
x=273